OpenStudy (anonymous):

A circle is defined by the equation x2 + y2 – x – 2y – = 0. What are the coordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

THE BLANK IS 11 OVER 4

OpenStudy (cwrw238):

convert the equation by completing squares x2 + y2 – x – 2y – 2.75 = 0 (x - 0.5)^2 - 0.25 + (y - 1)^2 -1- 2.75 = 0 (x - 0.5)^2 + (y - 1)^2 = 4

OpenStudy (cwrw238):

i dont like fractions so i changed them to decimals

OpenStudy (cwrw238):

now compare this last equation to the general equation of the circle and you'll get your answer (x - a)^2 + (y - b)^2 = r^2 here (a,b) is the center and r = radius

OpenStudy (cwrw238):

do you follow what I've written?

OpenStudy (anonymous):

YES THANKS, YOU HELPED A LOT. GOD BLESS YOU :-)

OpenStudy (cwrw238):

thank you - yw

OpenStudy (cwrw238):

if u need more help just post

OpenStudy (anonymous):

OK THANKS. HAVE A BLESS DAY